OpenStudy (anonymous):

Can I get some help with trigonometric identities? Specifically the power-reducing identity.

OpenStudy (anonymous):

Rewrite \[\tan ^{4}x sinx\]in terms of containing only first powers of sine and cosine.

zepdrix (zepdrix):

\[\tan^4x=\frac{ \sin^4x }{ \cos^4x }\] Hmm where can go from there :)

OpenStudy (anonymous):

Thanks. I can work it from there. I hadn't thought of doing it that way.